# Dressed basis sets for the modeling of exchange interactions in double quantum dots

### v1

@@ V Time-dependent simulations
As an illustration, we compute the spin funnel of the singlet-triplet qubit based on the squeezed DQD of the previous section [59, 60]. The electronic structure of this qubit at magnetic field $B_{x}=100$ mT is plotted as a function of the detuning voltage in Fig. 12a. The $S(1,1)$ anti-crosses the $S(0,2)$ state at $\delta V_{\mathrm{d}}=\delta V_{\mathrm{d}}^{*}=8.37$ mV. We start from the $S(0,2)$ state at $\delta V_{\mathrm{d}}>\delta V_{\mathrm{d}}^{*}$, then pulse diabatically (rise time $\tau_{r}=2$ ns) to some finite detuning $\delta V_{\mathrm{d}}<\delta V_{\mathrm{d}}^{*}$ in order to initialize the qubit in a singlet state. We next wait for $\tau_{w}=65$ ns and pulse diabatically back to the measurement point $\delta V_{\mathrm{d}}=9.74$ mV where the triplets are blocked. We finally measure the probability $P_{S}$ to return to the $S(0,2)$ state after this sequence. The time-dependent Schrödinger equation is solved on a grid of times $t=t_{0},\,...,\,t_{N}$; for that purpose, the evolution operator $U(t_{n+1},\,t_{n})=\exp[i\hat{H}(t_{n+1/2})(t_{n}-t_{n+1})/\hbar]$ is computed in each time interval $[t_{n},\,t_{n+1}]$ from the exact diagonalization of the Hamiltonian $\hat{H}(t_{n+1/2})$ in the dressed $M^{\prime}=112$ basis set [with $t_{n+1/2}=(t_{n}+t_{n+1})/2$]. We do not account for decoherence and relaxation in this calculation.
@@
Figure 12: (a) Electronic structure of the same squeezed DQD as in Fig. 11, as a function of the detuning voltage $\delta V_{\mathrm{d}}$ ($V_{\mathrm{J}}=-19.8$ meV, $B=100$ mT). The dashed vertical line indicates the position of the measurement point in the time-dependent simulations. The inset is a close-up on the $ST_{-}$ anti-crossing. (b) Funnel pattern computed in the squeezed DQD. The magnetic field is oriented along $\bm{x}$ in both panels. The top scales are the detuning energy with respect to the $S(1,1)/S(0,2)$ anti-crossing.
@@
The singlet return propability $P_{S}$ is plotted as a function of detuning and magnetic field amplitude in Fig. 12b. The magnetic field is oriented along $\bm{x}$. This figure exhibits the characteristic funnel structure concurring with the $ST_{-}$ anti-crossing. At these points of the $(\delta V_{d},B)$ plane, the $ST_{-}$ mixing is strong and the initial singlet state undergoes rapid $ST_{-}$ oscillations leading to a drop in $P_{S}$. The width of the funnel lines is, therefore, proportional to the singlet-triplet gap $\Delta_{ST_{-}}$. Along these lines $P_{S}$ is also modulated by the total, $\propto\tau_{w}$ phase accumulated during the $ST_{-}$ oscillations.
@@
There is a second set of lines clearly visible outside the funnel. They correspond to $ST_{0}$ oscillations at large enough $\delta\varepsilon$ and/or magnetic field. Indeed, the $S$ and $T_{0}$ states get mixed by the imbalance of Zeeman splittings $\Delta E_{Z}\propto(g_{L}^{*}-g_{R}^{*})B$ that competes with the exchange energy. This gives rise to a modulation of the return probability essentially periodic with the magnetic field deep in the $(1,1)$ charge state.
@@
Figure 13: (a) Singlet return probability $P_{S}(\tau_{w})$ as a function of the wait time $\tau_{w}$ at $\delta V_{\mathrm{d}}=0$ ($B=25$ mT). (b) Fourier transform ${\hat{P}}_{S}(\nu)$ of $P_{S}(\tau_{w})$. The vertical dashed line is the expected frequency of the $ST_{0}$ oscillations, $\nu=31.08$ MHz, and the vertical dotted line, the expected frequency of the $\lvert\Downarrow\Uparrow\rangle/T_{-}$ oscillations, $\nu=37.5$ MHz.
[…]
@@
Such maps and time traces are experimentally used to characterize the electronic structure of singlet-triplet qubits and demonstrate the existence of detuning-dependent $ST_{0}$ oscillations needed for their manipulation [59]. The dressed basis set provides a very accurate description of the dependence of the singlet and triplet states on detuning while allowing for efficient time-dependent calculations. It achieves, therefore, an excellent balance between accuracy and efficiency. The optimization of the operation point of this singlet-triplet qubit, as well as decoherence go beyond the scope of this paper, and will be addressed in future works.
@@ VI Conclusions
We have investigated the exchange interactions in double quantum dots with an efficient computational method. For that purpose, we have built a dressed basis set of two-particle wave functions that captures the main correlations that build up in full CI method. This dressed basis set can reproduce the low-energy singlet and triplet states over the whole operational gate voltages range with as few as a hundred basis functions, about fifty times less than in the CI. This speeds up intensive many-body calculations such as time-dependent simulations. We have applied this methodology to a double hole quantum dot in a germanium heterostructure. We have discussed the physics of the exchange interactions in this double dot, and we have highlighted the prominent role of Coulomb correlations. We have also analyzed the interplay between confinement, strains and Rashba interactions at finite magnetic field. We find that the spin-flip tunneling terms responsible for anisotropic exchange and singlet-triplet mixings are essentially driven by the imbalance of $g$-factors in this system, the inter-dot Rashba interactions being small even in the presence of inhomogeneous strains. This opens the way to the modeling of complex operations in Loss-DiVincenzo, singlet-triplet and exchange-only qubits with realistic layouts, disorder and noise.
@@ Acknowledgements
This work was supported by the “France 2030” program (PEPR PRESQUILE-ANR-22-PETQ-0002) and by the Horizon Europe Framework Program (grant agreement 101174557 QLSI2).
Appendix A Implementation details
A.1 Convergence and time-reversal symmetry
The calculated exchange interactions (e.g., Fig. 4) may become noisy when the barrier is closed ($V_{\mathrm{J}}\approx 0$ V) and the exchange energy gets in the neV range. Actually, all quantities (single-particle energies, $U_{ijkl}$’s…) must be converged with a better-than-neV accuracy to address this range, which is very demanding (and close to the computational limits, the spectrum of the FD hamiltonian spanning tens of eVs). It is also important not to break time-reversal symmetry (${\cal T}$) in the CI calculations, to avoid spurious mixings between singlets and triplets resulting in meaningless exchange splittings. To prevent (small) symmetry breakings by numerical inaccuracies, we always add $\lvert\psi_{1}\rangle$ and ${\cal T}\lvert\psi_{1}\rangle$ instead of both members $\lvert\psi_{1}\rangle$ and $\lvert\psi_{2}\rangle$ of a given Kramers pair to the single-particle basis set, and re-enforce a posteriori the time-reversal symmetry relations on the $W_{ijkl}$’s (namely, average the elements expected equal or conjugate by permutation of the indices).
@@ A.1 Convergence and time-reversal symmetry
A.2 Gauge invariance
Strict gauge invariance can hardly be enforced in a finite basis set as the latter can not accommodate arbitrary phase variations if the vector potential $\bm{A}$ gets large. Practically, we did not, however, experience significant differences with usual gauges (symmetric $\bm{A}=\bm{B}\times\bm{r}/2$, Landau…) centered on the DQD. Moreover, the model remains gauge-invariant in the linear response regime when both the matrices $N_{\mathrm{\alpha\beta}}$ and the elements of the $M_{\mathrm{\alpha}}$’s coupling different Kramers pairs can be neglected [50]. This is actually the regime of validity of the $g$-matrix formalism, which holds in the present cases [43]. We did, therefore, neglect the $N_{\mathrm{\alpha\beta}}$’s in the all calculations.

### v2

@@ V Time-dependent simulations
As an illustration, we compute the spin funnel of the singlet-triplet qubit based on the squeezed DQD of the previous section [59, 60]. The electronic structure of this qubit at magnetic field $B_{x}=100$ mT is plotted as a function of the detuning voltage in Fig. 12a. The $S(1,1)$ anti-crosses the $S(0,2)$ state at $\delta V_{\mathrm{d}}=\delta V_{\mathrm{d}}^{*}=8.37$ mV. We start from the $S(0,2)$ state at $\delta V_{\mathrm{d}}>\delta V_{\mathrm{d}}^{*}$, then pulse diabatically (rise time $\tau_{r}=2$ ns) to some smaller detuning $\delta V_{\mathrm{d}}=\delta V_{w}<\delta V_{\mathrm{d}}^{*}$ in order to initialize the qubit in a singlet state. We next wait for $\tau_{w}=65$ ns and pulse diabatically back to the measurement point $\delta V_{\mathrm{d}}=9.74$ mV where the triplets are blocked. We finally measure the probability $P_{S}$ to return to the $S(0,2)$ state after this sequence. The time-dependent Schrödinger equation is solved on a grid of times $t=t_{0},\,...,\,t_{N}$; for that purpose, the evolution operator $U(t_{n+1},\,t_{n})=\exp[i\hat{H}(t_{n+1/2})(t_{n}-t_{n+1})/\hbar]$ is computed in each time interval $t_{n},\,t_{n+1}]$ from the exact diagonalization of the Hamiltonian $\hat{H}(t_{n+1/2})$ [with $t_{n+1/2}=(t_{n}+t_{n+1})/2$]. This allows for large time steps at the wait point where the Hamiltonian is constant. $\hat{H}(t)$ can either be computed in the dressed ($M^{\prime}=112$) or full CI ($M=4560$) basis sets. We do not account for decoherence and relaxation in this calculation.
@@
Figure 12: (a) Electronic structure of the same squeezed DQD as in Fig. 11, as a function of the detuning voltage $\delta V_{\mathrm{d}}$ ($V_{\mathrm{J}}=-19.8$ meV, $B=100$ mT). The dashed vertical line indicates the position of the measurement point in the time-dependent simulations. The inset is a close-up on the $ST_{-}$ anti-crossing. (b) Funnel pattern computed in the squeezed DQD, as a function of the detuning voltage $\delta V_{w}$ at the wait point and magnetic field amplitude $B$. The magnetic field is oriented along $\bm{x}$ in both panels and the Hamiltonians are computed in the dressed $M^{\prime}=112$ basis set. The top scales are the detuning energy with respect to the $S(1,1)/S(0,2)$ anti-crossing.
@@
The singlet return probability $P_{S}$ is plotted as a function of $\delta V_{w}$ and $B$ in Fig. 12b. The magnetic field is oriented along $\bm{x}$. This figure, computed in the dressed basis set, exhibits the characteristic funnel structure concurring with the $ST_{-}$ anti-crossing. At these points of the $(\delta V_{w},B)$ plane, the $ST_{-}$ mixing is strong and the initial singlet state undergoes rapid $ST_{-}$ oscillations leading to a drop in $P_{S}$. The width of the funnel lines is, therefore, proportional to the singlet-triplet gap $\Delta_{ST_{-}}$. Along these lines $P_{S}$ is also modulated by the total, $\propto\tau_{w}$ phase accumulated during the $ST_{-}$ oscillations.
@@
There is a second set of lines clearly visible outside the funnel. They correspond to $ST_{0}$ oscillations at large enough magnetic field and detuning energy $\delta\varepsilon_{w}$ at the wait point. Indeed, the $S$ and $T_{0}$ states get mixed by the imbalance of Zeeman splittings $\Delta E_{Z}\propto(g_{L}^{*}-g_{R}^{*})B$ that competes with the exchange energy. This gives rise to a modulation of the return probability essentially periodic with the magnetic field deep in the $(1,1)$ charge state.
@@
Figure 13: (a) Singlet return probability $P_{S}(\tau_{w})$ as a function of the wait time $\tau_{w}$ at $\delta V_{w}=0$ ($B=25$ mT). The line is computed in the dressed basis set, and the dots in the full CI basis set. (b) Fourier transform ${\hat{P}}_{S}(\nu)$ of $P_{S}(\tau_{w})$ in the dressed basis set. The vertical dashed line is the expected frequency of the $ST_{0}$ oscillations, $\nu=31.08$ MHz, and the vertical dotted line, the expected frequency of the $\lvert\Downarrow\Uparrow\rangle/T_{-}$ oscillations, $\nu=37.5$ MHz.
Additionally, we plot in Fig. 13 the return probability $P_{S}(\tau_{w})$ as a function of the wait time $\tau_{w}$ at $\delta V_{w}=0$, and its Fourier transform ${\hat{P}}_{S}(\nu)$. The magnetic field is $B_{x}=25$ mT. At this operation point, the qubit shall primarily undergo $ST_{0}$ oscillations. Fig. 13 highlights different features. First, the $ST_{0}$ oscillations are incomplete ($P_{S}$ does not reach zero) due to the finite exchange at zero detuning. Indeed, $S$ and $T_{0}$ are not fully split into $\lvert\Uparrow\Downarrow\rangle$ and $\lvert\Downarrow\Uparrow\rangle$ states, so that an initial singlet state does not rotate around an axis of the equatorial plane of the $ST_{0}$ Bloch sphere that would bring it through the $T_{0}$ state [59]. Second, there is a clear beating between a main oscillation at $\nu=31.08$ MHz and a secondary oscillation at $\nu=37.5$ MHz. The former is the expected $ST_{0}$ oscillation, while the frequency of the latter matches the $\lvert\Downarrow\Uparrow\rangle/T_{-}$ splitting. This results from a small residual $ST_{-}$ mixing, and from non fully diabatic passages through the nearby $ST_{-}$ anti-crossing, which give rise to state preparation and measurement errors $P_{S}(0)<1$].
@@
As illustrated in Fig. 13a, the time-dependent calculations in the dressed basis set are in very good agreement with those in the full CI basis set, but are 55 000 times faster [^11] We emphasize, though, that the cost of the method we have chosen to solve the time-dependent Schrodinger equation scales as $M^{3}$ (with $M$ the dimension of the basis set). Other methods (Trotter, …) may show a better ($\propto M^{2}$) scaling. Anyhow, the gain in the dressed basis set remains very significant.. The dressed basis achieves, therefore, an excellent balance between accuracy and efficiency. Such maps and time traces are experimentally used to characterize the electronic structure of singlet-triplet qubits and demonstrate the existence of detuning-dependent $ST_{0}$ oscillations needed for their manipulation [59]. The optimization of the operation point of this singlet-triplet qubit, as well as decoherence go beyond the scope of this paper, and will be addressed in future works.
@@ VI Conclusions
We have investigated the exchange interactions in double quantum dots with an efficient computational method. For that purpose, we have built a dressed basis set of two-particle wave functions that captures the main correlations that build up in full CI method. This dressed basis set can reproduce the low-energy singlet and triplet states over the whole operational gate voltages range with as few as a hundred basis functions, about fifty times less than in the CI. This speeds up intensive many-body calculations such as time-dependent simulations. We have applied this methodology to a double hole quantum dot in a germanium heterostructure. We have discussed the physics of the exchange interactions in this double dot, and we have highlighted the prominent role of Coulomb correlations. We have also analyzed the interplay between confinement, strains and Rashba interactions at finite magnetic field. We find that the spin-flip tunneling terms responsible for anisotropic exchange and singlet-triplet mixings are essentially driven by the imbalance of $g$-factors in this system, the inter-dot Rashba interactions being small even in the presence of inhomogeneous strains. The methodology of this work can be extended to at least three dots (but at an increasing cost for the CI calculations needed to set-up the dressed basis set), and shall also capture long-range spin-dependent dipole interactions in the presence of strong spin-orbit coupling [141, 142, 143]. This opens the way to the modeling of complex operations in Loss-DiVincenzo, singlet-triplet and exchange-only qubits with realistic layouts, disorder and noise.
@@ Acknowledgements
This work was supported by the “France 2030” program (PEPR PRESQUILE-ANR-22-PETQ-0002), by the French National Research Agency (project InGeQT), and by the Horizon Europe Framework Program (grant agreement 101174557 QLSI2).
Appendix A Luttinger-Kohn Hamiltonian
The HH and LH envelope functions, $\varphi_{n}^{\nu}$ fulfill a set of differential equations defined by the Luttinger-Kohn Hamiltonian:
$H_{\mathrm{LK}}(\{V_{\mathrm{G}}^{0}\},\bm{B})=H_{\mathrm{K}}+H_{\varepsilon}+H_{\mathrm{Z}}+V_{t}(\bm{r})\mathbb{1}_{4}\,$, (30)
where $H_{\mathrm{K}}$ is the kinetic energy, $H_{\varepsilon}$ describes the effects of strains, $H_{\mathrm{Z}}$ is the Zeeman Hamiltonian and $\mathbb{1}_{4}$ is the $4\times 4$ identity matrix. $H_{\mathrm{K}}$ and $H_{\varepsilon}$ share the same generic form in the $J_{z}=\{+\tfrac{3}{2},+\tfrac{1}{2},-\tfrac{1}{2},-\tfrac{3}{2}\}$ basis set:
$H_{\mathrm{K/\varepsilon}}=\begin{pmatrix}P+Q&-S&R&0\\ -S^{\dagger}&P-Q&0&R\\ R^{\dagger}&0&P-Q&S\\ 0&R^{\dagger}&S^{\dagger}&P+Q\end{pmatrix}\,$, (31)
where, for $H_{\mathrm{K}}$,
$P_{\mathrm{K}}$ $=\frac{1}{2m_{0}}\gamma_{1}(p_{x}^{2}+p_{y}^{2}+p_{z}^{2})$ (32a)
$Q_{\mathrm{K}}$ $=\frac{1}{2m_{0}}\gamma_{2}(p_{x}^{2}+p_{y}^{2}-2p_{z}^{2})$ (32b)
$R_{\mathrm{K}}$ $=\frac{1}{2m_{0}}\sqrt{3}\left[-\gamma_{2}(p_{x}^{2}-p_{y}^{2})+2i\gamma_{3}\{p_{x},\,p_{y}\}\right]$ (32c)
$S_{\mathrm{K}}$ $=\frac{1}{2m_{0}}2\sqrt{3}\gamma_{3}\{p_{x}-ip_{y},\,p_{z}\}\,$, (32d)
with $\{A,\,B\}=\tfrac{1}{2}(AB+BA)$, and, for $H_{\varepsilon}$,
$P_{\varepsilon}$ $=-a_{v}(\varepsilon_{xx}+\varepsilon_{yy}+\varepsilon_{zz})$ (33a)
$Q_{\varepsilon}$ $=-\frac{1}{2}b_{v}(\varepsilon_{xx}+\varepsilon_{yy}-2\varepsilon_{zz})$ (33b)
$R_{\varepsilon}$ $=\frac{\sqrt{3}}{2}b_{v}(\varepsilon_{xx}-\varepsilon_{yy})-id_{v}\varepsilon_{xy}$ (33c)
$S_{\varepsilon}$ $=-d_{v}(\varepsilon_{xz}-i\varepsilon_{yz})\,.$ (33d)
Here $\mathbf{p}$ is the momentum, $m_{0}$ is the free electron mass, and $\gamma_{1}$, $\gamma_{2}$, $\gamma_{3}$ are the Luttinger parameters that characterize the hole masses. The $\varepsilon_{\alpha\beta}$ are the strains; $a_{v}$ is the hydrostatic, $b_{v}$ the uniaxial and $d_{v}$ the shear deformation potential of the valence band. The form of Eq. (31), which couples different $J_{z}$’s through the $R$ and $S$ terms, embodies the action of SOC in the valence band. The Zeeman Hamiltonian $H_{\mathrm{Z}}=2\mu_{B}(\kappa\mathbf{B}\cdot\mathbf{J}+q\mathbf{B}\cdot\mathbf{J}^{3})$ describes the action of the magnetic field on the Bloch functions, with $\mathbf{J}$ the spin $\tfrac{3}{2}$ operator, $\mathbf{J}^{3}\equiv(J_{x}^{3},J_{y}^{3},J_{z}^{3})$, $\mu_{B}$ the Bohr magneton, and $\kappa$, $q$ the isotropic and cubic Zeeman parameters. The action of $\mathbf{B}$ on the envelopes of the hole is accounted for by the substitution $\mathbf{p}\to-i\hbar\bm{\nabla}+e\mathbf{A}$ in $H_{\mathrm{K}}$, with $\mathbf{A}=\frac{1}{2}\mathbf{B}\times\mathbf{r}$ the magnetic vector potential.
Appendix B Implementation details
B.1 Convergence and time-reversal symmetry
The calculated exchange interactions (e.g., Fig. 4) may become noisy when the barrier is closed ($V_{\mathrm{J}}\approx 0$ V) and the exchange energy gets in the neV range. Actually, all quantities (single-particle energies, $U_{ijkl}$’s, etc.) must be converged with a better-than-neV accuracy to address this range, which is very demanding (and close to the computational limits, the spectrum of the FD hamiltonian spanning tens of eVs). It is also important not to break time-reversal symmetry ($\cal T$}) in the CI calculations, to avoid spurious mixings between singlets and triplets resulting in meaningless exchange splittings. To prevent (small) symmetry breakings by numerical inaccuracies, we always add $\lvert\psi_{1}\rangle$ and ${\cal T}\lvert\psi_{1}\rangle$ instead of both members $\lvert\psi_{1}\rangle$ and $\lvert\psi_{2}\rangle$ of a given Kramers pair to the single-particle basis set, and re-enforce a posteriori the time-reversal symmetry relations on the $W_{ijkl}$’s (namely, average the elements expected equal or conjugate by permutation of the indices).
@@ B.1 Convergence and time-reversal symmetry
B.2 Gauge invariance
Strict gauge invariance can hardly be enforced in a finite basis set as the latter can not accommodate arbitrary phase variations if the vector potential $\bm{A}$ gets large. Practically, we did not, however, experience significant differences with usual gauges (symmetric $\bm{A}=\bm{B}\times\bm{r}/2$, Landau, etc…) centered on the DQD. Moreover, the model is gauge-invariant in the linear response regime where the matrices $N_{\mathrm{\alpha\beta}}$ can be discarded, if the elements of the $M_{\mathrm{\alpha}}$’s coupling different Kramers pairs of $H_{0}$ can also be neglected [50]. The linear response (the $g$-matrix formalism) holds in the present cases [43]. We did, therefore, neglect the $N_{\mathrm{\alpha\beta}}$’s in the all calculations.